OpenStudy (domebotnos):

A system of equations is shown below: 6x – 5y = 5 3x + 5y = 4 The x-coordinate of the solution to this system of equations is _____. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (compassionate):

Hi, so if you're trying to solve for x then you need to eliminate y. You alreadysee -5y and + 5y will cancel, so then simply add and solve or x.

OpenStudy (domebotnos):

how would that be? or look like?

OpenStudy (compassionate):

Okay, so you have 6x – 5y = 5 3x + 5y = 4 Your system is already set up for elimination. So you can add the whole systems together and your y terms will eliminate each other 6x + 3x - 5y + 5y = 5 + 4 9x = 9 This is because -5y + 5y = 0. Now we can solve for x.

OpenStudy (domebotnos):

im still confused...

OpenStudy (domebotnos):

...

OpenStudy (compassionate):

Which part are you confused about?